### v1

# Volumes, Lorenz-like Templates, and Braids

Abstract.
Lorenz-like templates are obtained from applying half twists to the two strips of the Lorenz template. Ghrist proved that infinitely many of these templates are universal, meaning they contain all links in the 3-sphere. In this paper, we generalise the bunch algorithm to construct links embedded in universal Lorenz-like templates and provide an upper volume bound that is quadratic in the trip number. We also establish a correspondence between links embedded in these universal Lorenz-like templates and generalised T–links. While generalised T-links have appeared in the knot theory literature before, we prove that they are, in fact, all links in the 3-sphere. Thus, we can regard each link in the 3-sphere as a generalised T-link, which has an associated Lorenz link. We use the upper bound obtained from the generalised bunch algorithm to provide an upper volume bound for all hyperbolic link complements in the 3-sphere. Such a bound grows quadratically in terms of the braid indexes of the associated Lorenz links.
1. Introduction
Meteorologist E. N. Lorenz, in his pursuit to understand weather dynamics, uncovered a system of three ordinary differential equations in $\mathbb{R}^{3}$ [20]:
$\frac{dx}{dt}=10(y-x),\quad\frac{dy}{dt}=28x-y-xz,\quad\frac{dz}{dt}=xy-\frac{8}{3}z.$
This system reveals complex dynamics, including numerous knotted closed periodic orbits, known as Lorenz links.
Guckenheimer and Williams [14], along with Tucker [22], established that Lorenz links can be investigated by examining links within the Lorenz template, an embedded branched surface in $\mathbb{R}^{3}$. They demonstrated that every Lorenz link can be continuously deformed to a link embedded in the Lorenz template, and every link embedded in the template is a Lorenz link. Thus, Lorenz links can be defined as links embedded in the Lorenz template.
This motivated Birman and Williams to define a more general concept, called templates, which are more general compact branched $2$-manifold that is equipped with a semiflow, see Section 2 for more information about templates. Furthermore, they demonstrated that the periodic orbits of any Anosov flow in a $3$-manifold can be embedded in a template [5].
From the Lorenz template, we can obtain infinitely many templates by applying half twists on both strips of the Lorenz templates. They are called Lorenz-like templates and denoted by $\mathscr{L}(u,v)$, where $u$ and $v$ represent the numbers of half twists applied to the left and right strips of the Lorenz template, respectively. Figure 1 illustrates the Lorenz-like template $\mathscr{L}(1,2)$. A link that can be embedded in $\mathscr{L}(u,v)$ is called a Lorenz-like link. We note that the templates $\mathscr{L}(u,v)$ and $\mathscr{L}(v,u)$ contain the same family of links.
A universal template in the 3-sphere is a template that has a representative of every ambient isotopy class of links in the 3-sphere. Ghrist proved in [12] that the Lorenz-like templates $\mathscr{L}(0,v)$ are universal if $v<0$. Therefore, the Lorenz-like templates $\mathscr{L}(0,v)$ with $v<0$ contain a diagram for each link in the 3-sphere.
In this paper, we study links in the $3$-sphere using the diagrams induced by the Lorenz-like templates $\mathscr{L}(u,v)$. We start by examining their hyperbolic volumes. This investigation leads us to derive an upper bound for the volumes of the (possibly empty) hyperbolic pieces of the complements of the Lorenz-like links in the template $\mathscr{L}(u,v)$, with $u$, $v$ being even integers, using the trip number. When a link is embedded in $\mathscr{L}(u,v)$, it is associated with a natural number known as the trip number. This number represents the number of strands in the Lorenz-like template that traverse from the left strip to the right strip. For a formal definition, refer to Section 2. This number was first defined by Birman and Williams for Lorenz links in the Lorenz template [3], which was later proved to be the braid indexes of the Lorenz links by Franks and Williams in [10]. We relate the volume and the trip number of link as follows.
Theorem 3.4.
Let $u$, $v$ be even integers. Consider $L$ a link in the $3$-sphere embedded in a Lorenz-like template $\mathscr{L}(u,v)$ with trip number equal to $\overline{n}$. Let $\mathrm{Vol}(L)$ denote the sum of the volumes of the hyperbolic pieces of $\mathbb{S}^{3}-L$. Then,
$\mathrm{Vol}(L)\leq 12\ v_{\textup{tet}}(\overline{n}^{2}+3\overline{n}),$
where $v_{\textup{tet}}\approx 1.01494$ is the volume of the regular ideal tetrahedron.
Since the templates $\mathscr{L}(0,v)$ with $v<0$ are universal, Theorem 3.4 provides an upper bound for the volumes of the hyperbolic pieces of all link complements in the $3$-sphere.
We then investigate a simplification of the diagrams derived from the Lorenz-like templates $\mathscr{L}(0,v)$ to diagrams that are closer to the ones we find in the knot theory literature. This was inspired by the equivalence between Lorenz links and T-links demonstrated by Birman and Kofman in [2]. This results in a simpler braid description for Lorenz-like links in terms of a generalisation of torus links or T-links, defined as follows.
Definition 4.11.
Let $r_{1},\dots,r_{k},r_{k+1},s_{1},\dots,s_{k},s_{k+1},d$ be integers such that $2\leq r_{1}<\dots<r_{k}<r_{k+1}$, $s_{1},\dots,s_{k}>0$, $s_{k+1}\in\mathbb{Z}-\{0\}$, and $d\geq 0$. If $s_{k+1}<0$ then the generalised T-link
$T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}),d)$
is defined to be the closure of the following braid
$(\sigma_{1+d}\dots\sigma_{d+r_{1}-1})^{s_{1}}\dots(\sigma_{1+d}\dots\sigma_{d+r_{k}-1})^{s_{k}}(\sigma_{d+r_{k+1}-1}^{-1}\sigma_{d+r_{k+1}-2}^{-1}\dots\sigma_{1+d}^{-1})^{s_{k+1}}$
otherwise
$(\sigma_{1+d}\dots\sigma_{d+r_{1}-1})^{s_{1}}\dots(\sigma_{1+d}\dots\sigma_{d+r_{k}-1})^{s_{k}}(\sigma_{1+d}\dots\sigma_{d+r_{k+1}-1})^{s_{k+1}}.$
Generalised T-links $T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}),d)$ differ from standard T-links because the second coordinate $s_{k+1}$ of the largest pair $(r_{k+1},s_{k+1})$ can be either positive or negative. Additionally, generalised T-links can be splittable links, which can occur, for example, when $d>0$. In contrast, T-links cannot be splittable links.
We prove that generalised T-links are in equivalence with Lorenz-like links in the Lorenz-like template $\mathscr{L}(0,2v)$ for some non-positive integer $v$.
Theorem 4.13.
Generalised T-links $T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}),d)$ with $s_{k+1}\geqslant vr_{k+1}$ are in equivalence with Lorenz-like links in the Lorenz-like template $\mathscr{L}(0,2v)$ if $v\leq 0$.
We combine Theorem 4.13 and the fact that the templates $\mathscr{L}(0,2v)$ are universal if $v<0$ to obtain a simple parameterization for all links in the 3-sphere given by the following theorem.
Corollary 4.16.
Generalised T-links consist of all links in $\mathbb{S}^{3}$.
We note that two generalised T-links with different set of parameters can represent the same link, see Theorem 4.8.
In particular, we note that this family of links has appeared in knot theory literature before. More precisely, Champanerkar, Futer, Kofman, Neumann, and Purcell found an upper bound for the sum of the volumes of the hyperbolic pieces of a family of links that they called T-links [6]. However, we conclude from Corollary 4.16 that the collection of all T-links in their definition includes all links in the 3-sphere, up to a finite number of unknots.
Corollary 4.17.
Any link in the 3-sphere is the union of a $T$-link in [6, Definition 1.2] and a finite number of unknots. The upper bound in [6, Theorem 1.7] applies to the sum of the volumes of the hyperbolic pieces of all links in the 3-sphere.
In [17, Example 5.7], Tetsuya defined a similar generalisation of T-links, which takes the form $T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}),0)$ in our definition of generalised T-links. However, he allowed all $s_{i}$ to be either positive or negative. Similarly, we conclude that any link in the 3-sphere can be expressed as the union of a generalised T-link defined in [17, Example 5.7] and a finite number of unknots.
We then combine Theorem 3.4 and Theorem 4.13 to obtain the following upper bound for the volumes of the hyperbolic pieces of all links in the 3-sphere.
Theorem 4.18.
Consider $L=T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}),d)$ a generalised T-link with $s_{k+1}\geqslant nr_{k+1}$ for some integer $n$. Let $\mathrm{Vol}(L)$ denote the sum of the volumes of the hyperbolic pieces of $\mathbb{S}^{3}-L$. Furthermore, let $\beta$ be the braid index of the Lorenz link
$L^{\prime}=T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}-nr_{k+1})).$
Then,
$\mathrm{Vol}(L)\leq 12v_{tet}(\beta^{2}+3\beta).$
Furthermore, this provides an upper bound for the volume of all hyperbolic pieces of each link complement in the 3-sphere.
See the end of Section 4 for a discussion of the upper bound for the volumes that we obtain from Theorem 4.18 and Champanerkar, Futer, Kofman, Neumann and Purcell [6].
Other upper bounds for the hyperbolic volumes of hyperbolic links in the 3-sphere have been found, for example, Lackenby, Agol, and Thurston found an upper bound in terms of the twist number [18], Adams obtained an upper bound in terms of the number of triple crossings [1]. However, they can be difficult to compare with Theorem 4.18, as there are questions related to minimality in some of these results. For example, it is not obvious how one can find the minimal number of twist regions of a link in [18], or the minimal braid index of the associated Lorenz links of a link in Theorem 4.18.
Finally, we apply the fundamental theorem of Wallace and Lickorish [24, 19], along with Thurston’s result on volume reduction under Dehn filling [21, Proposition 6.5.2], to establish the first upper bound for the volumes of the hyperbolic pieces of all closed, orientable 3-manifolds.
Theorem 4.19.
For any closed, orientable and connected $3$-manifold $M$, there exists a generalised T-link
$L\coloneqq T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}),d)$
with $s_{k+1}\geqslant nr_{k+1}$ for some integer $n$ such that $M$ can be obtained by Dehn filling the complement of $L$ in the $3$-sphere. Furthermore, let $\mathrm{Vol}(M)$ be the sum of the volumes of the hyperbolic pieces of $M$ and let $\beta$ be the braid index of the Lorenz link
$T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}-nr_{k+1})).$
We have
$\mathrm{Vol}(M)\leq 12v_{tet}(\beta^{2}+3\beta).$
We observe that the upper bound in [6, Theorem 1.7] can also be applied to the sum of the volumes of the hyperbolic pieces of all closed, orientable 3-manifolds when we consider Corollary 4.17.
1.1. Organisation
In section 2, we introduce the generalised bunch algorithm for constructing links embedded in Lorenz-like templates, this provides the basics for proving Theorem 3.4 in Section 3. In Section 4, we prove Theorem 4.13 along with the other remaining theorems. To prove Theorem 4.13, we utilize a slightly different proof than that used by Birman and Kofman in [2] to demonstrate the equivalence of Lorenz links and T-links. The remaining theorems are derived from Theorem 3.4 and Theorem 4.13.
1.2. Acknowledgements
We thank Tali Pinsky, Jessica Purcell, and Pierre Dehornoy for helpful conversations. The first author was supported by the Australian Research Council, grant DP210103136.
2. The generalised bunch algorithm and Lorenz-like templates
A template is a compact branched $2$-manifold with nonempty boundary. Such a branched surface is endowed with a semiflow constructed from the joining and splitting charts as shown in [13, Figure 2.4]. A famous example is the Lorenz template (see Figure 4, left), which is a geometric model introduced by Guckenheimer and Williams [15, 25] for studying Lorenz flow. Tucker [23] later justified the geometric model. Lorenz links can thus be considered as links embedded in the Lorenz template and studied using symbolic dynamics (see [25, 4]).
In [16], the second and third authors developed the bunch algorithm for constructing Lorenz links in the Lorenz template when a code word is given. The bunch algorithm provided the basics and insights for finding an upper volume bound for Lorenz link complements that is independent of the word exponents and quadratic in the trip number, which is the braid index of the Lorenz link [11, 4, 2]. It turns out that the bunch algorithm can be generalised to infinitely many universal templates, which are templates embedded in the $3$-sphere that contain every link in the $3$-sphere up to ambient isotopies (see also [13, Definition 3.2.15]). The following provides more details.
Definition 2.1.
Let $u$, $v$ be integers. The $(u,v)$-Lorenz-like template $\mathscr{L}(u,v)$ is a template obtained by cutting the Lorenz template along the branch line, glue the left strip back with $u$ half twists, and glue the right strip back with $v$ half twists. Figure 1 shows an example with positive half twists. Negative half twists are obtained by switching the crossings.
We call a link embedded in $\mathscr{L}(u,v)$ a $(u,v)$-Lorenz-like link, or simply a Lorenz-like link.
$x$
$y$
Figure 1. Lorenz-like template $\mathscr{L}(1,2)$.
Proposition 2.2 ([12] and Proposition 3.2.16 in [13]).
If $v$ is a negative integer, then the Lorenz-like template $\mathscr{L}(0,v)$ is universal.
For example, Proposition 2.2 implies that every link in the $3$-sphere is a Lorenz-like link embedded in the template $\mathscr{L}(0,-2)$.
Similar to code words of modular links (see [4] and [16, Definition 3.1 (4)] for more details), which are Lorenz links, we have code words and labelled code words for Lorenz-like links using the symbols $x$ and $y$ as shown in Figure 1.
From here on, we use the terminologies introduced in [16] to state the more generalised bunch algorithm and the corresponding machinery for us to show the upper bound results in later sections.
Let $L$ be a Lorenz-like link with $c$ link components. We denote the corresponding $c$ labelled code words for $L$ by the following:
(*) $\begin{dcases}w_{1}&=x_{1}^{k_{1}}y_{1}^{l_{1}}x_{2}^{k_{2}}y_{2}^{l_{21}}\ldots x_{n_{1}}^{k_{n_{1}}}y_{n_{1}}^{l_{n_{1}}},\\ w_{2}&=x_{n_{1}+1}^{k_{n_{1}+1}}y_{n_{1}+1}^{l_{n_{1}+1}}x_{n_{1}+2}^{k_{n_{1}+2}}y_{n_{1}+2}^{l_{n_{1}+2}}\ldots x_{n_{1}+n_{2}}^{k_{n_{1}+n_{2}}}y_{n_{1}+n_{2}}^{l_{n_{1}+n_{2}}},\\ &\ldots\\ w_{c}&=x_{\Sigma(c)+1}^{k_{\Sigma(c)+1}}y_{\Sigma(c)+1}^{l_{\Sigma(c)+1}}x_{\Sigma(c)+2}^{k_{\Sigma(c)+2}}y_{\Sigma(c)+2}^{l_{\Sigma(c)+2}}\ldots x_{\overline{n}}^{k_{\overline{n}}}y_{\overline{n}}^{l_{\overline{n}}}\end{dcases}$
The notation $\Sigma(c)$ represents the sum $\Sigma_{i=1}^{c-1}n_{i}$ of the word periods for words $w_{1},\ldots,w_{c-1}$ if $c>1$. The symbol $\overline{n}$ denotes the sum of all word periods for $L$, that is, $\overline{n}=\Sigma(c+1)=\Sigma_{i=1}^{c}n_{i}$. Note that we use the term trip number and sum of word periods interchangeably.
Theorem 2.3 (Bunches in Lorenz-like links).
Let $u$, $v$ be even integers. Each $(u,v)$-Lorenz-like link $L$ with labelled code words in (*2) is ambient isotopic to a union of bunches of turns that are ordered from left to right in the split template according to the order of turns listed in each bunch, and the union of bunches satisfy both conditions below:
  1. (1)
    For any $i\in[1,\overline{n}]\cap\mathbb{Z}$, the $x_{i}^{k_{i}}$-arc starts from a point in the $-{k_{i}}^{\textup{th}}$ interval.
  2. (2)
    For any $j\in[1,\overline{n}]\cap\mathbb{Z}$, the $y_{j}^{l_{j}}$-arc starts from a point in the $+{l_{j}}^{\textup{th}}$ interval.
Proof.
Since $u$, $v$ are even integers, the $(u,v)$-Lorenz-like template is obtained by cutting along the branch line of the Lorenz template, and making $\frac{u}{2}$ full twists on the left strip and $\frac{v}{2}$ full twists on the right strip before gluing back. As the number of full twists are integer-valued, we can get a split template that is the same as that of the Lorenz template (see example in [16, Figure 2]), and the corresponding sets of code words are the same for both templates. The statement then follows from the proof of [16, Theorem 3.8]. ∎
The concept order within full bunches in Lorenz link introduced in [16] can apply to Lorenz-like links. Readers may refer to [16, Section 3.2] for more details and examples of finding order within full bunches given the code words of a Lorenz-like link.
Following from Theorem 2.3 and [16, Proposition 3.9], we have the following generalised bunch algorithm for constructing the Lorenz-like link in $\mathscr{L}(u,v)$ (where $u$, $v$ are even integers) whenever code words are given:
Corollary 2.4 (The generalised bunch algorithm).
Let $u$, $v$ be even integers. Let $L$ be a $(u,v)$-Lorenz-like link embedded in $\mathscr{L}(u,v)$ with the sum of all word periods denoted by $\overline{n}$. Suppose $K$ is a link component of $L$ with labelled code word $w=x_{1}^{k_{1}}y_{1}^{l_{1}}\ldots x_{n}^{k_{n}}y_{n}^{l_{n}}$. The link component $K$ can be constructed in the split template using the following steps:
  1. (1)
    Mark the branch line with integers $-k_{\mu},\ldots,+l_{\lambda}$, where $k_{\mu}$ and $l_{\lambda}$ are a maximal $x$- and $y$-exponents among all words associated to $L$ respectively.
  2. (2)
    Determine the orders within the full bunches in the Lorenz link $L$ using [16, Proposition 3.9]. In each negative (or positive resp.) interval, draw $\overline{n}$ points with $x$-base (or $y$-base resp.) labels from left to right according to the order within the full $x$-bunch (or full $y$-bunch resp.).
  3. (3)
    Construct $K$ by first drawing the $x_{1}^{k_{1}}$-arc, which starts from the $x_{1}$-point in the $-{k_{1}}^{\textup{th}}$ interval and ends at the $y_{1}$-point in the $+{l_{1}}^{\textup{th}}$ interval. Then, draw the $y_{1}^{l_{1}}$-arc which starts from the $y_{1}$-point in the $+{l_{1}}^{\textup{th}}$ interval and ends at the $x_{2}$-point in the $-{k_{2}}^{\textup{th}}$ interval. Continue drawing the arcs for subwords $x_{2}^{k_{2}},y_{2}^{l_{2}},\ldots,x_{n}^{k_{n}},y_{n}^{l_{n}}$ with known starting and ending points to obtain $K$ in the split template.
By repeating Step (3) for each link component of $L$ and making $u$ and $v$ half twists on the left and right strips of the split template respectively, the Lorenz-like link $L$ embedded in the Lorenz-like template can be obtained.
3. An upper volume bound for all links in the $3$-sphere
In this section, we provide an upper bound for the sum of the volumes of the hyperbolic pieces of all hyperbolic link complements in the $3$-sphere using the techniques developed in Section 2 and [16]. Such an upper bound is quadratic in the minimal trip number.
One of the results we use for finding the upper volume bound is Theorem 1.5 in [7], it can be stated as follows:
Theorem 3.1 (Theorem 1.5 in [7]).
 [^1] Abuse of notation: The symbol $\mathcal{L}$ in Theorem 1 is considered as an embedding in the function composition and as a set otherwise.
Let $M$ be a Seifert-fibred space over a hyperbolic $2$-orbifold $B$ with bundle projection map $\mathcal{P}\colon\thinspace M\to B$. If $\mathcal{L}$ is a link embedded in $M$ such that $\mathcal{P}\circ\mathcal{L}$ is a collection of loops that intersect themselves only transversely and finitely many times with exactly two pre-image points for each self-intersection point, then the simplicial volume of $M{\smallsetminus}\mathcal{L}$ satisfies the following inequality:
$\norm{M{\smallsetminus}\mathcal{L}}\leq 8\ \iota(\mathcal{P}(\mathcal{L}),\mathcal{P}(\mathcal{L})),$
where $\iota(\mathcal{P}(\mathcal{L}),\mathcal{P}(\mathcal{L}))$ is the number of self-intersections of $\mathcal{P}(\mathcal{L})$.
For details on simplicial volume and its relationship with hyperbolic volume, in particular the fact that the number of tetrahedra in any topological ideal triangulation of a finite volume cusped hyperbolic manifold is an upper bound for its volume, readers may refer to [21, 6.5.2], [21, 6.5.4] and [21, 6.1.7].
Definition 3.2 (Parent manifold of Lorenz-like link).
[…]
  • the parent link $L_{\textup{p}}$ (see [16, Definition 4.5] for its construction),
  • the unknots $V_{a_{1}},\ldots,V_{a_{\#l}}$, where each $V_{j}$ encircles all the blue vertical line segment(s) (as defined in Step (3) of [16, Definition 4.5]) that is/are about to intersect the annulus $A^{y}_{j}$, and
  • the unknots $V_{L}$ and $V_{R}$ that encircle the left and right sets of annuli respectively.
See Figure 2 for example.
$V_{2}$
$V_{3}$
$V_{6}$
$U$
[…]
Proposition 3.3.
[…]
Proof.
The argument is similar to the argument for [16, Proposition 4.8] as the generalised bunch algorithm (Corollary 2.4) holds. The differences between the parent manifold in [16, Proposition 4.8] and the parent manifold here are:
  1. (1)
    The former leaves the trefoil knot component not Dehn-filled while the later does not drill out the trefoil knot component.
  2. (2)
    The former does not have twists on the two strips of the Lorenz template while the later may have full twists, which require extra Dehn fillings on $V_{L}$ and $V_{R}$.
Hence, other than the Dehn fillings mentioned in the argument of [16, Proposition 4.8], we apply $(\frac{1}{u/2})$-Dehn filling on $V_{L}$ and apply $(\frac{1}{v/2})$-Dehn filling on $V_{R}$. The desired complement of the Lorenz-like link $L$ can thus be obtained. ∎
Theorem 3.4.
Let $u$, $v$ be even integers. Consider $L$ a link in the $3$-sphere embedded in a Lorenz-like template $\mathscr{L}(u,v)$ with trip number equal to $\overline{n}$. Let $\mathrm{Vol}(L)$ denote the sum of the volumes of the hyperbolic pieces of $\mathbb{S}^{3}-L$. Then,
$\mathrm{Vol}(L)\leq 12\ v_{\textup{tet}}(\overline{n}^{2}+3\overline{n}),$
where $v_{\textup{tet}}\approx 1.01494$ is the volume of the regular ideal tetrahedron.
Proof.
[…]
By [16, Lemma 4.10], the parent link $L_{p}$ can be continuously deformed to a location such that the number of self-intersections of $\mathcal{P}(L_{p})$ is at most $\frac{3}{2}\overline{n}(\overline{n}-1)$.
The unknot(s) $V_{a_{1}},\ldots,V_{a_{\#l}}$ for untwisting unwanted full twist(s) contribute at most $2\overline{n}$ self-crossings of $\mathcal{P}(\mathcal{L}_{L})$ because there are a total of $\overline{n}$ (blue) vertical strand(s) approaching the $y$-annuli and the number of crossings that each unknot creates is twice the number of strands it encircles. (See Figure 3 for example.)
Hence, the total number of self-crossings in $\mathcal{P}(\mathcal{L}_{L}{\smallsetminus}(V_{L}\cup V_{R}))$ is
$\frac{3}{2}\overline{n}(\overline{n}-1)+2\overline{n}=\frac{3}{2}\overline{n}^{2}+\frac{1}{2}\overline{n}.$
With the addition of the two loops $V_{L}$ and $V_{R}$ in the parent manifold, the total number of self-crossings in $\mathcal{P}(\mathcal{L}_{L})$ becomes
$\frac{3}{2}\overline{n}^{2}+\frac{1}{2}\overline{n}+2*2*\overline{n}=\frac{3}{2}\overline{n}^{2}+\frac{9}{2}\overline{n}$
By [7, Theorem 1.5], we have
[…]
Since Dehn fillings do not increase volumes [21, Proposition 6.5.2], it follows that
[…]
an upper bound that is quadratic in terms of the sum of word period(s). ∎
$V_{2}$
$V_{3}$
$V_{6}$
$V_{R}$
$V_{L}$
$U$
Figure 3. The projection of the link $\mathcal{L}_{K}=K_{\textup{p}}\cup V_{L}\cup V_{R}\cup V_{2}\cup V_{3}\cup V_{6}$ in the $14$-punctured disc bounded by the unknot $U$. The green loops $V_{L}$ and $V_{R}$ produce $2*2*5=20$ more crossings in the link projection.
Note that a link can be embedded in infinitely many universal templates and the trip number of the same link may differ in different templates (see Theorem 4.8 for a related result).
Proposition 2.2 states that the Lorenz-like template $\mathscr{L}(0,v)$ is universal for any negative integer $v$, and the generalised bunch algorithm works for all these universal templates with $v$ being an even integer. Motivated from the above observations, we define the following invariant of link:
Definition 3.5 (Minimal trip number).
Let $v$ be an even integer. For any link $L$ in the $3$-sphere, the minimal trip number of $L$ is the smallest integer $\overline{n}$ such that the following holds: There exists an even negative integer $v$ such that $L$ is embedded in the Lorenz-like template $\mathscr{L}(0,v)$ and $\overline{n}$ is the trip number of the embedded $L$ in $\mathscr{L}(0,v)$.
The following corollary follows from Theorem 3.4.
Corollary 3.6.
If $L$ is a link in the $3$-sphere with minimal trip number $\tilde{n}$, then the following inequality holds:
$\mathrm{Vol}(L)\leq 12\ v_{\textup{tet}}(\tilde{n}^{2}+3\tilde{n}),$
where $v_{\textup{tet}}\approx 1.01494$ is the volume of the regular ideal tetrahedron.
4. A new twist on Lorenz-like links
In this section, we provide a simpler braid description for all Lorenz-like links that can be embedded into the Lorenz-like templates $\mathscr{L}(0,v)$ and $\mathscr{L}(u,0)$, as obtained by Birman and Kofman for the Lorenz template in [2]. As a consequence, we obtain a simpler braid description for all links in the 3-sphere. This simpler braid description can be viewed as a generalisation of torus links and provides a simple parametrization for all links in the 3-sphere. Unexpectedly, Champanerkar, Futer, Kofman, Neumann, and Purcell have previously worked with this family of links in [6, Definition 1.2], but they didn’t realize that they are, in fact, all links in the 3-sphere, as we will demonstrate.
$x$
$y$
Figure 4. Left: Lorenz template. Right: Braid Lorenz template.
The Lorenz template $\mathscr{L}(0,0)$ (see Definition 2.1) produces a braid template for the Lorenz links. This is achieved by cutting and opening the Lorenz template, as illustrated in Figure 4.
A Lorenz braid in the braid template is formed by groups of overcrossing and undercrossing strands, with the overcrossing strands located among the leftmost strands and the undercrossing strands among the rightmost strands at the top of the Lorenz braid. Furthermore, the overcrossing strands never cross each other, and the undercrossing strands never cross each other. A Lorenz braid is characterized by the permutation associated with its strands.
If a Lorenz braid has no overcrossing strands, or equivalently, if the Lorenz link has a trip number equal to zero, then the Lorenz link is the unlink with trivial components that are parallel to the two boundary components of the Lorenz template. Therefore, we consider that the Lorenz braid has at least one overcrossing strand. Furthermore, given a link $L$ in the Lorenz template, we can obtain infinitely many splittable links by adding trivial loops that are parallel to the left boundary component of the template $\mathscr{L}(0,0)$. For now, when we consider a link in the Lorenz template, we disregard these trivial loops that are parallel to the left boundary component.
A Lorenz braid is well-defined by the permutation associated with the overcrossing strands because once the positions of the overcrossing strands are set, there is only one way to place the undercrossing strands since they do not cross. The overcrossing strands are inclined from the left to the right. If an overcrossing strand begins at $i$, then it ends at $(i+r_{i})$ with $r_{i}>0$. This provides the following sequence of natural numbers:
$1\leq r_{1}\leq r_{2}\leq\dots\leq r_{p-1}\leq r_{p}.$
We can assume that $r_{1}>1$ because if $r_{1}=1$, we can apply a destabilization move on the left of this Lorenz braid of $L$ to obtain a new Lorenz braid for $L$ with $p-1$ overcrossing strands. Hence, we can assume that $r_{1}\geq 2$. Furthermore, since it is possible that $r_{i}=r_{i+1}$, if we associate the strands of the same slope into the same group, we can rewrite this sequence of natural numbers as the following vector, called a Lorenz vector,
$\vec{d}_{L}=\langle d_{1}^{s_{1}},\dots,d_{k}^{s_{k}}\rangle,$
with $2\leq d_{1}<\dots<d_{k}$, $\{r_{1},r_{2},\dots,r_{p}\}=\{d_{1},\dots,d_{k}\}$, and $s_{1}+\dots+s_{k}=p$.
Figure 5. The Lorenz vector $\vec{d}_{L}=\langle 2^{2},4^{2},7^{3},8^{2}\rangle$.
Let $\vec{d}_{L}=\langle d_{1}^{s_{1}},\dots,d_{k}^{s_{k}}\rangle$ be a Lorenz vector as in Figure 4. Denote by $(d_{1},s_{1})$ the braid
$(\sigma_{1}\dots\sigma_{d_{1}})(\sigma_{2}\dots\sigma_{d_{1}+1})\dots(\sigma_{s_{1}}\dots\sigma_{d_{1}+s_{1}-1}).$
More generally, for $i>1$, denote by $(d_{i},s_{i})$ the braid
$(\sigma_{s_{1}+\dots+s_{i-1}+1}\dots\sigma_{s_{1}+\dots+s_{i-1}+d_{i}})\dots(\sigma_{s_{1}+\dots+s_{i-1}+s_{i}}\dots\sigma_{s_{1}+\dots+s_{i-1}+s_{i}+d_{i}-1}).$
The Lorenz braid associated with the Lorenz vector $\vec{d}_{L}=\langle d_{1}^{s_{1}},\dots,d_{k}^{s_{k}}\rangle$ is the following braid
$(d_{k},s_{k})(d_{k-1},s_{k-1})\dots(d_{2},s_{2})(d_{1},s_{1}),$
which has $s_{1}+\dots+s_{k}+d_{k}$ strands.
Let $a\geq b>1$. We denote by $\Delta_{b,a}$ the braid
$(\sigma_{a-b+1}\sigma_{a-b+2}\dots\sigma_{a-1})(\sigma_{a-b+1}\sigma_{a-b+2}\dots\sigma_{a-2})\dots(\sigma_{a-b+1}\sigma_{a-b+2})(\sigma_{a-b+1}).$
In other words, $\Delta_{b,a}$ denotes a positive half twist on the last $b$ strands of the trivial braid with $a$ strands. On the other hand, we denote by $\Delta_{b,a}^{-1}$ a negative half twist on the last $b$ strands of the trivial braid with $a$ strands, as shown in the following braid
$(\sigma_{a-1}^{-1}\sigma_{a-2}^{-1}\dots\sigma_{a-b+2}^{-1}\sigma_{a-b+1}^{-1})(\sigma_{a-1}^{-1}\sigma_{a-2}^{-1}\dots\sigma_{a-b+2}^{-1})\dots(\sigma_{a-1}^{-1}\sigma_{a-2}^{-1})(\sigma_{a-1}^{-1}).$
Every link $L$ in the Lorenz-like template $\mathscr{L}(0,n)$ is obtained from a Lorenz link $L^{\prime}$, called the Lorenz link associated with $L$, by applying $n$ half twists on the strands of $L^{\prime}$ in the right strip of the Lorenz template $\mathscr{L}(0,0)$. If the Lorenz braid of $L^{\prime}$ has no overcrossing strands, or equivalently, if the Lorenz link $L^{\prime}$ has a trip number equal to zero, then $L$ is a collection of trivial loops that are parallel to the two boundary components of the template with $n$ half twists applied to some of them.
Furthermore, given a link $L$ in the Lorenz-like template $\mathscr{L}(0,n)$, we can obtain infinitely many splittable links by adding trivial loops that are parallel to the left boundary component of the template $\mathscr{L}(0,n)$. For now, when we consider a link in the Lorenz-like template $\mathscr{L}(0,n)$, we disregard these trivial loops that are parallel to the left boundary component.
Moreover, if the Lorenz braid of $L^{\prime}$ has only one overcrossing strand, or equivalently, if the Lorenz link $L^{\prime}$ has a trip number equal to one, and there is only one strand in the right strip of the template $\mathscr{L}(0,0)$, then $L$ and $L^{\prime}$ are both the same unlink. We consider then that there is at least one overcrossing strand and at least two strands in the right strip of the template $\mathscr{L}(0,0)$ in the Lorenz braid of $L^{\prime}$.
Let $\vec{d}_{L^{\prime}}=\langle d_{1}^{s_{1}},\dots,d_{k}^{s_{k}}\rangle$ be the Lorenz vector of $L^{\prime}$, called the Lorenz vector associated with the link $L$. Then, there exists a natural number $d\geq 0$ such that $L$ is the closure of the following braid with $s_{1}+\dots+s_{k}+d_{k}+d$ strands
$(d_{k},s_{k})(d_{k-1},s_{k-1})\dots(d_{2},s_{2})(d_{1},s_{1})\Delta^{n}_{d_{k}+d,s_{1}+\dots+s_{k}+d_{k}+d.}$
The number $d$ can be greater than zero if we add trivial loops that are parallel to the boundary component of the strip of the Lorenz template that is twisted to obtain the Lorenz-like template $\mathscr{L}(0,n)$. We observe that the trip number of $L$ in the Lorenz-like template $\mathscr{L}(0,n)$ is equal to the trip number of $L^{\prime}$ in the Lorenz template $\mathscr{L}(0,0)$ because the $n$ half twists applied along the strands of $L^{\prime}$ in the right strip of $\mathscr{L}(0,n)$ do not change the trip number.
Proposition 4.1.
Let $\vec{d}_{L^{\prime}}=\langle d_{1}^{s_{1}},\dots,d_{k}^{s_{k}}\rangle$ be a Lorenz vector with $k>1$ or if $k=1$ then $s_{1}>1$. Consider $\beta=\sigma_{i_{1}}\dots\sigma_{i_{a}}$ a braid with more than or equal to $s_{1}+\dots+s_{k}+d_{k}$ strands and no crossings $\sigma_{i_{j}}$ with $i_{j}\leq s_{1}+\dots+s_{k}$. Then, there is an isotopy that takes the braid
$B^{\prime}=(d_{k},s_{k})(d_{k-1},s_{k-1})\dots(d_{2},s_{2})(d_{1},s_{1})\beta$
to the braid
$B^{\prime\prime}=(d_{k},s_{k})\dots(d_{2},s_{2})(d_{1},s_{1}-1)\beta^{\prime}(\sigma_{s_{1}+\dots+s_{k}}\dots\sigma_{s_{1}+\dots+s_{k}+d_{1}-2})$
with $\beta^{\prime}=\sigma_{i_{1}-1}\dots\sigma_{i_{a}-1}$, where $B^{\prime\prime}$ has one less strand than $B^{\prime}$.
Proof.
The $(d_{1}+1)$-th strand at the bottom of the sub-braid $(d_{k},s_{k})\dots(d_{1},s_{1}-1)$ anticlockwise goes around the braid closure once to end in the $(1)$-st strand at the bottom. Thus, the $(d_{1}+1)$-th strand is connected to the $(1)$-st strand by an over strand that goes one time around the braid closure as illustrated by the red strand in the second drawing of Figure 6.
$\beta$
$\beta$
$\beta^{\prime}$
$\beta^{\prime}$
Figure 6. This series of pictures (from left to right) illustrates how one strand from the closure of the leftmost braid can be reduced.
We push this over strand up and shrink it to reduce one strand from the braid $B$, as shown in the third drawing of Figure 6. After that, this strand becomes an over strand between the $(1)$-th and the $(d_{1})$-st strand yielding the braid
$(d_{k},s_{k})(d_{k-1},s_{k-1})\dots(d_{2},s_{2})(d_{1},s_{1}-1)(\sigma_{1}\dots\sigma_{d_{1}-1})\beta^{\prime},$
with $\beta^{\prime}=\sigma_{i_{1}-1}\dots\sigma_{i_{a}-1}.$
Since the first $s_{1}+\dots+s_{k}-1$ strands of the last braid are still overcrossing strands with the last crossings $\sigma_{i}$ with $i\geq d_{1}$, we can push the sub-braid $(\sigma_{1}\dots\sigma_{d_{1}-1})$ up to be on the top of the braid so that we obtain the braid
$(\sigma_{s_{1}+\dots+s_{k}}\dots\sigma_{s_{1}+\dots+s_{k}+d_{1}-2})(d_{k},s_{k})(d_{k-1},s_{k-1})\dots(d_{2},s_{2})(d_{1},s_{1}-1)\beta^{\prime}.$
Finally we push the sub-braid $(\sigma_{s_{1}+\dots+s_{k}}\dots\sigma_{s_{1}+\dots+s_{k}+d_{1}-2})$ anticlockwise around the braid closure to place it below the sub-braid $\beta^{\prime}$ as illustrated in the last drawing of Figure 6. After that, we obtain the braid
$B^{\prime\prime}=(d_{k},s_{k})\dots(d_{2},s_{2})(d_{1},s_{1}-1)\beta^{\prime}(\sigma_{s_{1}+\dots+s_{k}}\dots\sigma_{s_{1}+\dots+s_{k}+d_{1}-2}).\qed$
Lemma 4.2.
Let $\vec{d}_{L^{\prime}}=\langle d_{1}^{s_{1}}\rangle$ be a Lorenz vector with $s_{1}\geq 1$ and $\beta=\sigma_{i_{1}}\dots\sigma_{i_{a}}$ be a braid with more than or equal to $s_{1}+d_{1}$ strands and no crossings $\sigma_{i_{j}}$ with $i_{j}\leq s_{1}$. Then, there is an isotopy that takes the braid
$B^{\prime}=(d_{1},s_{1})\beta$
to the braid
$B^{\prime\prime}=\beta^{\prime}(\sigma_{1}\dots\sigma_{d_{1}-1})^{s_{1}}$
with $\beta^{\prime}=\sigma_{i_{1}-s_{1}}\dots\sigma_{i_{a}-s_{1}}$.
Proof.
After applying $s_{1}-1$ times Proposition 4.1 to $B^{\prime}$, we obtain the following braid, called $B_{s_{1}-1}$,
$(d_{1},1)\beta_{s_{1}-1}(\sigma_{2}\dots\sigma_{d_{1}})^{s_{1}-1}$
with $\beta_{s_{1}-1}=\sigma_{i_{1}-(s_{1}-1)}\dots\sigma_{i_{a}-(s_{1}-1)}.$
The sub-braid $\beta_{s_{1}-1}(\sigma_{2}\dots\sigma_{d_{1}})^{s_{1}-1}$ has no crossing $\sigma_{i_{j}}$ with $i_{j}=1$. Hence we can trivially destabilize $B_{s_{1}-1}$ on its left side to yield the braid
$(\sigma_{1}\dots\sigma_{d_{1}-1})\beta_{s_{1}}(\sigma_{1}\dots\sigma_{d_{1}-1})^{s_{1}-1}$
with $\beta_{s_{1}}=\sigma_{i_{1}-s_{1}}\dots\sigma_{i_{a}-s_{1}}.$ Then, we push the sub-braid $(\sigma_{1}\dots\sigma_{d_{1}-1})$ anticlockwise one time around the braid closure to obtain the braid
$\beta_{s_{1}}(\sigma_{1}\dots\sigma_{d_{1}-1})^{s_{1}}.$
So, we set $\beta^{\prime}=\beta_{s_{1}}$. ∎
Proposition 4.3.
Let $\vec{d}_{L^{\prime}}=\langle d_{1}^{s_{1}},\dots,d_{k}^{s_{k}}\rangle$ be a Lorenz vector and $\beta=\sigma_{i_{1}}\dots\sigma_{i_{a}}$ be a braid with more than or equal to $s_{1}+\dots+s_{k}+d_{k}$ strands and no crossings $\sigma_{i_{j}}$ with $i_{j}\leq s_{1}+\dots+s_{k}$. Then, there is an isotopy that takes the braid
$B^{\prime}=(d_{k},s_{k})(d_{k-1},s_{k-1})\dots(d_{1},s_{1})\beta$
to the braid
$(\sigma_{1}\dots\sigma_{d_{1}-1})^{s_{1}}(\sigma_{1}\dots\sigma_{d_{2}-1})^{s_{2}}\dots(\sigma_{1}\dots\sigma_{d_{k}-1})^{s_{k}}\beta^{\prime}.$
with $\beta^{\prime}=\sigma_{i_{1}-(s_{1}+\dots+s_{k})}\dots\sigma_{i_{a}-(s_{1}+\dots+s_{k})}$.
Proof.
If $k=1$, then $B^{\prime}$ is equivalent to the braid $B^{\prime\prime}=\beta^{\prime}(\sigma_{1}\dots\sigma_{d_{1}-1})^{s_{1}}$ with $\beta^{\prime}=\sigma_{i_{1}-s_{1}}\dots\sigma_{i_{a}-s_{1}}$ by Lemma 4.2. We push the sub-braid $(\sigma_{1}\dots\sigma_{d_{1}-1})^{s_{1}}$ clockwise one time around the braid closure to obtain the braid
$(\sigma_{1}\dots\sigma_{d_{1}-1})^{s_{1}}\beta^{\prime}.$
Consider then that $k>1$. We apply $s_{1}+\dots+s_{k-1}$ times Proposition 4.1 to $B^{\prime}$ to obtain the following braid
$B_{s_{1}+\dots+s_{k-1}}=(d_{k},s_{k})\beta_{s_{1}+\dots+s_{k-1}}(\sigma_{s_{k}+1}\dots\sigma_{s_{k}+d_{1}-1})^{s_{1}}$
$(\sigma_{s_{k}+1}\dots\sigma_{s_{k}+d_{2}-1})^{s_{2}}\dots(\sigma_{s_{k}+1}\dots\sigma_{s_{k}+d_{k-1}-1})^{s_{k-1}},$
with $\beta_{s_{1}+\dots+s_{k-1}}=\sigma_{i_{1}-(s_{1}+\dots+s_{k-1})}\dots\sigma_{i_{a}-(s_{1}+\dots+s_{k-1})}.$ Then, we apply Lemma 4.2 to $B_{s_{1}+\dots+s_{k-1}}$ to obtain the following braid
$\beta_{s_{1}+\dots+s_{k}}(\sigma_{1}\dots\sigma_{d_{1}-1})^{s_{1}}(\sigma_{1}\dots\sigma_{d_{2}-1})^{s_{2}}\dots(\sigma_{1}\dots\sigma_{d_{k}-1})^{s_{k}},$
with $\beta_{s_{1}+\dots+s_{k}}=\sigma_{i_{1}-(s_{1}+\dots+s_{k})}\dots\sigma_{i_{a}-(s_{1}+\dots+s_{k})}.$ Finally, we push the sub-braid $(\sigma_{1}\dots\sigma_{d_{1}-1})^{s_{1}}\dots(\sigma_{1}\dots\sigma_{d_{k}-1})^{s_{k}}$ clockwise one time around the braid closure to obtain the braid
$(\sigma_{1}\dots\sigma_{d_{1}-1})^{s_{1}}(\sigma_{1}\dots\sigma_{d_{2}-1})^{s_{2}}\dots(\sigma_{1}\dots\sigma_{d_{k}-1})^{s_{k}}\beta_{s_{1}+\dots+s_{k}}.\qed$
For every integer $n$ we have a family of links defined as follows:
Definition 4.4.
Let $r_{1},\dots,r_{k},r_{k+1},s_{1},\dots,s_{k},d$ be non-negative integers. The $T^{n}$-link
$T^{n}((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1};d))$
is defined to be the closure of the following braid
$(\sigma_{1+d}\sigma_{2+d}\dots\sigma_{d+r_{1}-1})^{s_{1}}\dots(\sigma_{1+d}\sigma_{2+d}\dots\sigma_{d+r_{k}-1})^{s_{k}}\Delta^{n}_{r_{k+1},d+r_{k+1}}$
if $k+1>1$, $2\leq r_{1}<\dots<r_{k}\leq r_{k+1}$, all $s_{i}>0$, and $d\geq 0$, or
$\Delta^{n}_{r_{k+1},d+r_{k+1}}$
if $k+1=1,r_{1}\geq 1$ and $d\geq 0$, where $T^{n}(1;d)$ is the unlink with $d+1$ components given by the trivial braid with $d+1$ strands.
T-links are defined as follows: for $2\leq r_{1}<\dots<r_{k}$, and all $s_{i}>0$, the T-link $T((r_{1},s_{1}),\dots,(r_{k},s_{k}))$ is defined to be the closure of the following braid
$(\sigma_{1}\sigma_{2}\dots\sigma_{r_{1}-1})^{s_{1}}(\sigma_{1}\sigma_{2}\dots\sigma_{r_{2}-1})^{s_{2}}\dots(\sigma_{1}\sigma_{2}\dots\sigma_{r_{k}-1})^{s_{k}}.$
Birman and Kofman showed that T-links are equivalent to Lorenz links [2].
If $n=0$, then $T^{0}$-links are the same as T-links, up to a number of unknots. This occurs because Birman and Kofman in [2] didn’t consider splittable links that have trivial components parallel to the left boundary component of the Lorenz template. This completely fine because the study of a splittable link is done by studying its components separately. However, we need to consider all components if we want to find all links in $\mathbb{S}^{3}$.
Theorem 4.5.
Lorenz-like links in the Lorenz-link template $\mathscr{L}(0,n)$ are equivalent to $T^{n}$-links. In other words, every link that can be embedded into the Lorenz-like template $\mathscr{L}(0,n)$ is a $T^{n}$-link, and every $T^{n}$-link can be embedded into the Lorenz-like template $\mathscr{L}(0,n)$.
Proof.
Consider $L$ a link in the Lorenz-link template $\mathscr{L}(0,n)$. Let $L^{\prime}$ be a Lorenz link associated with the link $L$.
If $L$ has a trip number equal to zero in the Lorenz-link template $\mathscr{L}(0,n)$, then $L^{\prime}$ is a collection of trivial loops that are parallel to the two boundary components of the Lorenz template $\mathscr{L}(0,0)$. Since $L$ is obtained from $L^{\prime}$ by applying $n$ half twists on the strands in the right strip of the Lorenz template $\mathscr{L}(0,0)$, then $L$ is a collection of trivial loops with $n$ half twists applied to some of them. Hence, $L$ is the $T^{n}$-link $T^{n}(r_{1};d)$ with $r_{1}\geq 1$ and $d\geq 0$.
If the Lorenz braid of $L^{\prime}$ has one overcrossing strand, or equivalently, if the Lorenz link $L^{\prime}$ has a trip number equal to one and only one strand in the right strip of the template $\mathscr{L}(0,0)$, then $L$ and $L^{\prime}$ are both the same unlink $T^{n}(1;d)$ with $d\geq 0$.
Consider now the case where there is at least one overcrossing strand and at least two strands in the right strip of the template $\mathscr{L}(0,0)$ in the Lorenz braid of $L^{\prime}$. The link $L$ can be an obvious splittable link by having trivial loops that are parallel to the left boundary component of the template $\mathscr{L}(0,n)$. If link $L$ has some of these trivial loops, we disregard them for now. Let $\vec{d}_{L^{\prime}}=\langle d_{1}^{s_{1}},\dots,d_{k}^{s_{k}}\rangle$ be the Lorenz vector associated with $L$, where $d_{1},\dots,d_{k},d,s_{1},\dots,s_{k}$ are natural numbers such that $2\leq d_{1}<\dots<d_{k}$ and $s_{i}>0$. Furthermore, there exists a natural number $d\geq 0$ such that $L$ is the closure of the following braid with $p=s_{1}+\dots+s_{k}+d_{k}+d$ strands
$(d_{k},s_{k})(d_{k-1},s_{k-1})\dots(d_{2},s_{2})(d_{1},s_{1})\Delta^{n}_{d_{k}+d,p}.$
By Proposition 4.3, there is an isotopy that takes the closure of the last braid to the closure of the braid
$(\sigma_{1}\dots\sigma_{d_{1}-1})^{s_{1}}\dots(\sigma_{1}\dots\sigma_{d_{k}-1})^{s_{k}}\Delta^{n}_{d_{k}+d,d_{k}+d},$
which represents the $T^{n}$-link $T^{n}((d_{1},s_{1}),\dots,(d_{k},s_{k}),(d_{k}+d;0))$.
If now we consider that $L$ has trivial loops that are parallel to the left boundary component of the template $\mathscr{L}(0,n)$, then there exists a number $d^{\prime}\geq 0$ such that $L$ is the closure of the braid
$(\sigma_{1+d^{\prime}}\dots\sigma_{d^{\prime}+d_{1}-1})^{s_{1}}\dots(\sigma_{1+d^{\prime}}\dots\sigma_{d^{\prime}+d_{k}-1})^{s_{k}}\Delta^{n}_{d_{k}+d,d_{k}+d+d^{\prime}},$
which represents the $T^{n}$-link $T^{n}((d_{1},s_{1}),\dots,(d_{k},s_{k}),(d_{k}+d;d^{\prime}))$.
Similarly, if we apply the reverse procedure, we can see that every $T^{n}$-link can be embedded into the Lorenz-like template $\mathscr{L}(0,n)$. ∎
In particular, we obtain [2, Theorem 1] as a specific case of Theorem 4.5 and simplify its proof.
Corollary 4.6.
Consider $d,d^{\prime},d_{1},\dots,d_{k},s_{1},\dots,s_{k}$ non-negative integers such that $2\leq d_{1}<\dots<d_{k}$ and $s_{i}>0$. The T-link $T((d_{1},s_{1}),\dots,(d_{k},s_{k}))$ is a Lorenz link associated with the $T^{n}$-link $T^{n}((d_{1},s_{1}),\dots,(d_{k},s_{k}),(d_{k}+d;d^{\prime}))$.
Proof.
It follows from the proof of Theorem 4.5
Corollary 4.7.
Lorenz-like links in the Lorenz-link template $\mathscr{L}(m,0)$ are in equivalence with $T^{m}$-links. In other words, every link that can be embedded into the Lorenz-like template $\mathscr{L}(m,0)$ is a $T^{m}$-link and every $T^{m}$-link can be embedded into the Lorenz-like template $\mathscr{L}(m,0)$
Proof.
It follows from Theorem 4.5 since the templates $\mathscr{L}(m,0)$ and $\mathscr{L}(0,m)$ have the same family of links. ∎
Theorem 4.8.
Every T-link has infinitely many representations with different trip numbers in each template $\mathscr{L}(0,2n)$ with $n<0$.
Proof.
Consider the T-link $L=T((r_{1},s_{1}),\dots,(r_{k},s_{k}))$. By definition, the parameters of $L$ satisfy $2\leq r_{1}<\dots<r_{k}$, $s_{1},\dots,s_{k}>0$. However, we can consider that $s_{k}>1$ otherwise the T-link representation of $L^{\prime}$ can be destabilized on its right to yield the T-link $T((r_{1},s_{1}),(r_{2},s_{2}),\dots,(r_{k-1},s_{k-1}+1)).$ So we assume that $s_{k}>1$.
The T-link $L$ is given by the braid
$(\sigma_{1}\dots\sigma_{r_{1}-1})^{s_{1}}(\sigma_{1}\dots\sigma_{r_{2}-1})^{s_{2}}\dots(\sigma_{1}\dots\sigma_{r_{k}-1})^{s_{k}},$
which is equivalent to the braid
$(\sigma_{1}\dots\sigma_{r_{1}-1})^{s_{1}}\dots(\sigma_{1}\dots\sigma_{r_{k}-1})^{s_{k}-1}(\sigma_{1}\dots\sigma_{r^{\prime}_{k}-1})^{1}$
or
$(\sigma_{1}\dots\sigma_{r_{1}-1})^{s_{1}}\dots(\sigma_{1}\dots\sigma_{r_{k}-1})^{s_{k}-1}(\sigma_{1}\dots\sigma_{r^{\prime}_{k}-1})^{1-nr^{\prime}_{k}}(\sigma_{1}\dots\sigma_{r^{\prime}_{k}-1})^{nr^{\prime}_{k}},$
which is the same as
$(\sigma_{1}\dots\sigma_{r_{1}-1})^{s_{1}}\dots(\sigma_{1}\dots\sigma_{r_{k}-1})^{s_{k}-1}(\sigma_{1}\dots\sigma_{r^{\prime}_{k}-1})^{1-nr^{\prime}_{k}}\Delta^{2n}_{r^{\prime}_{k},r^{\prime}_{k}}$
for any $r^{\prime}_{k}\geq r_{k}$ and $n<0$. The last braid represents the $T^{2n}$-link
$L^{\prime}=T^{2n}((r_{1},s_{1}),\dots,(r_{k},s_{k}-1),(r^{\prime}_{k},1-nr^{\prime}_{k}),(r^{\prime}_{k};0)).$
Hence the T-link $L$ is equivalent to the $T^{2n}$-link $L^{\prime}$.
By Corollary 4.6,
$L^{\prime\prime}=T((r_{1},s_{1}),\dots,(r_{k},s_{k}-1),(r^{\prime}_{k},1-nr^{\prime}_{k}))$
is a Lorenz link associated with the Lorenz-like link $L^{\prime}$. The trip number of $L^{\prime}$ in the Lorenz-like template $\mathscr{L}(0,2n)$ is the same as the trip number of $L^{\prime\prime}$ in the Lorenz template $\mathscr{L}(0,0)$. But the trip number of a Lorenz link is equal to its braid index. This fact was conjectured by Birman and Williams in [3] and proved by Franks and Williams in [10]. The braid index of $L^{\prime\prime}$ is equal to $r^{\prime}_{k}$ since it has a positive braid with $r_{k}$ strands and a full twist on $r_{k}$ strands. This follows from the fact that a positive braid with a full twist reaches the braid index, as proved in [10, Corollary 2.4].
Therefore, the T-link $L=T((r_{1},s_{1}),\dots,(r_{k},s_{k}))$ has infinitely many representations of the form
$L^{\prime}=T^{2n}((r_{1},s_{1}),\dots,(r_{k},s_{k}-1),(r^{\prime}_{k},1-nr^{\prime}_{k}),(r^{\prime}_{k};0))$
in the template $\mathscr{L}(0,2n)$ with trip numbers equal to $r^{\prime}_{k}$ for any $r^{\prime}_{k}\geq r_{k}$ and $n<0$. We conclude that every T-link has infinitely many representations with different trip numbers in the template $\mathscr{L}(0,2n)$ for each $n<0$. ∎
Corollary 4.9.
Consider
$L=T^{2n}((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1};d))$
a $T^{2n}$-link in $\mathbb{S}^{3}$ with $n$ being any integer. Let $\mathrm{Vol}(L)$ denote the sum of the volumes of the hyperbolic pieces of $\mathbb{S}^{3}-L$. If $\beta$ is the braid index of the Lorenz link $T((r_{1},s_{1}),\dots,(r_{k},s_{k}))$, then
$\mathrm{Vol}(L)\leq 12v_{tet}(\beta^{2}+3\beta).$
Proof.
By Theorems 4.5 and 3.4, we have $\mathrm{Vol}(L)\leq 12v_{tet}(\tau^{2}+3\tau),$ where $\tau$ is the trip number of $L$ in the Lorenz-like template $\mathscr{L}(0,2n)$ By Corollary 4.6, $L^{\prime}=T((r_{1},s_{1}),\dots,(r_{k},s_{k}))$ is the Lorenz link associated with the Lorenz-like link $L$. The trip number of $L$ in the Lorenz-like temaplte $\mathscr{L}(0,2n)$ is the same as the trip number of $L^{\prime}$ in the Lorenz template $\mathscr{L}(0,0)$. But the trip number of a Lorenz link is equal to its braid index. Therefore,
$\mathrm{Vol}(L)\leq 12v_{tet}(\beta^{2}+3\beta),$
where $\beta$ is the braid index of the Lorenz link $T((r_{1},s_{1}),\dots,(r_{k},s_{k}))$. ∎
Corollary 4.10.
Consider $L=T((r_{1},s_{1}),\dots,(r_{k},s_{k}))$ a T-link in $\mathbb{S}^{3}$ with $s_{k}\geq r_{k}$ and $\lambda$ a positive integer such that $0\leq s_{k}-\lambda r_{k}<r_{k}$. Let $\mathrm{Vol}(L)$ denote the sum of the volumes of the hyperbolic pieces of $\mathbb{S}^{3}-L$. Furthermore, let $\beta$ be the braid index of the Lorenz link $T((r_{1},s_{1}),\dots,(r_{k-1},s_{k-1}),(r_{k},s_{k}-\lambda r_{k}))$. Then,
$\mathrm{Vol}(L)\leq 12v_{tet}(\beta^{2}+3\beta)<12v_{tet}(r_{k}^{2}+3r_{k}).$
Proof.
The Lorenz link
$L=T((r_{1},s_{1}),\dots,(r_{k-1},s_{k-1}),(r_{k},s_{k}-\lambda r_{k}),(r_{k},\lambda r_{k}))$
is the $T^{2\lambda}$-link
$L=T^{2\lambda}((r_{1},s_{1}),\dots,(r_{k-1},s_{k-1}),(r_{k},s_{k}-\lambda r_{k}),(r_{k};0))$
with associated Lorenz link equal to $L^{\prime}=T((r_{1},s_{1}),\dots,(r_{k-1},s_{k-1}),(r_{k},s_{k}-\lambda r_{k}))$ by Corollary 4.6. Then, by Corollary 4.9,
$\mathrm{Vol}(L)\leq 12v_{tet}(\beta^{2}+3\beta),$
where $\beta$ is the braid index of the Lorenz link $L^{\prime}$. Finally, we have that $\beta<r_{k}$ from [2, Corollary 8], [8, Section 3], or [9, Theorem 3.9]. ∎
Definition 4.11.
Let $r_{1},\dots,r_{k},r_{k+1},s_{1},\dots,s_{k},s_{k+1},d$ be integers such that $2\leq r_{1}<\dots<r_{k}<r_{k+1}$, $s_{1},\dots,s_{k}>0$, $s_{k+1}\in\mathbb{Z}-\{0\}$, and $d\geq 0$. The generalised T-link
$T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}),d)$
is defined to be the closure of the following braid
$(\sigma_{1+d}\dots\sigma_{d+r_{1}-1})^{s_{1}}\dots(\sigma_{1+d}\dots\sigma_{d+r_{k}-1})^{s_{k}}(\sigma_{1+d}\dots\sigma_{d+r_{k+1}-1})^{s_{k+1}}.$
If $s_{k+1}<0$, then
$(\sigma_{1+d}\dots\sigma_{d+r_{k+1}-2}\sigma_{d+r_{k+1}-1})^{s_{k+1}}=(\sigma_{d+r_{k+1}-1}^{-1}\sigma_{d+r_{k+1}-2}^{-1}\dots\sigma_{1+d}^{-1})^{-s_{k+1}}$
in the braid group. So, more concretely, in case $s_{k+1}<0$, then the generalised T-link $T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}),d)$ is the closure of the braid
$(\sigma_{1+d}\dots\sigma_{d+r_{1}-1})^{s_{1}}\dots(\sigma_{1+d}\dots\sigma_{d+r_{k}-1})^{s_{k}}(\sigma_{d+r_{k+1}-1}^{-1}\dots\sigma_{1+d}^{-1})^{-s_{k+1}}.$
Theorem 4.12.
Generalised T-links $T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}),d)$ with $s_{k+1}\geqslant nr_{k+1}$ and $T^{2n}$-links are in equivalence if $n\leq 0$. If $n>0$, then generalised T-links $T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}),d)$ with $s_{k+1}\geqslant nr_{k+1}$ together with the trivial links $T((2,1),d)$ are in equivalence with $T^{2n}$-links.
Proof.
The trivial links of the form $T((2,1),d)$ are $T^{2n}$-links of the form $T^{2n}(1;d)$. Now let $L$ be a generalised T-link $L=T((r_{1},s_{1}),\dots,(r_{k+1},s_{k+1}),d)$ with $s_{k+1}\geqslant nr_{k+1}$. Hence $s_{k+1}-nr_{k+1}\geq 0$. The link $L$ is given by the braid
$(\sigma_{1+d}\dots\sigma_{d+r_{1}-1})^{s_{1}}\dots(\sigma_{1+d}\dots\sigma_{d+r_{k}-1})^{s_{k}}(\sigma_{1+d}\dots\sigma_{d+r_{k+1}-1})^{s_{k+1}},$
which is equivalent to the braid
$(\sigma_{1+d}\dots\sigma_{d+r_{1}-1})^{s_{1}}\dots(\sigma_{1+d}\dots\sigma_{d+r_{k}-1})^{s_{k}}$
$(\sigma_{1+d}\dots\sigma_{d+r_{k+1}-1})^{s_{k+1}-nr_{k+1}}\Delta^{2n}_{r_{k+1},d+r_{k+1}},$
which represents the $T^{2n}$-link
$T^{2n}((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}-nr_{k+1}),(r_{k+1};d))$
(it becomes $T^{2n}((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1};d))$ in case $s_{k+1}-nr_{k+1}=0$).
Now let $L$ be a $T^{2n}$-link
$T^{2n}((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1};d))$
given by the braid
$(\sigma_{1+d}\sigma_{2+d}\dots\sigma_{d+r_{1}-1})^{s_{1}}\dots(\sigma_{1+d}\sigma_{2+d}\dots\sigma_{d+r_{k}-1})^{s_{k}}\Delta^{2n}_{r_{k+1},d+r_{k+1}},$
which is the same as the braid
$(\sigma_{1+d}\dots\sigma_{d+r_{1}-1})^{s_{1}}\dots(\sigma_{1+d}\dots\sigma_{d+r_{k}-1})^{s_{k}}(\sigma_{1+d}\sigma_{2+d}\dots\sigma_{d+r_{k+1}-1})^{nr_{k+1}}.$
Consider $r_{k}=r_{k+1}$. Then, this braid represents the generalised T-link
$T((r_{1},s_{1}),\dots,(r_{k-1},s_{k-1}),(r_{k},s_{k}+nr_{k}),d)$
if $s_{k}+nr_{k}\neq 0$, with $s_{k}+nr_{k}>nr_{k}$, otherwise
$T((r_{1},s_{1}),\dots,(r_{k-1},s_{k-1}),d),$
with $s_{k-1}>nr_{k-1}$ since $n$ is negative in this case. If $r_{k}\neq r_{k+1}$, the last braid represents the generalised T-link
$T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},nr_{k+1}),d)$
with $nr_{k+1}=nr_{k+1}.$
Finally consider a $T^{2n}$-link of the form $T^{2n}(r_{1};d)$. The link $L$ is given by the braid
$\Delta^{2n}_{r_{1},d+r_{1}}$
with $r_{1}\geq 1$ and $d\geq 0$, where $T^{2n}(1;d)$ is the unlink with $d+1$ components. So, if $r_{1}=1$, then $L$ is the generalised T-link $T((2,1),d)$, where we have $1>2n$ if $n\leq 0$. If $r_{1}>1$, then $L$ is the generalised T-link
$T((r_{1},nr_{1}),d),$
with $nr_{1}=nr_{1}$.
Therefore, if $n\leq 0$, then $T^{2n}$-links and generalised T-links
$T((r_{1},s_{1}),\dots,(r_{k+1},s_{k+1}),d)$
with $s_{k+1}\geqslant nr_{k+1}$ are equivalent. If $n>0$, then generalised T-links $T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}),d)$ with $s_{k+1}\geqslant nr_{k+1}$ together with the trivial links $T((2,1),d)$ are in equivalence with $T^{2n}$-links. ∎
Theorem 4.13.
Generalised T-links $T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}),d)$ with $s_{k+1}\geqslant vr_{k+1}$ are in equivalence with Lorenz-like links in the Lorenz-like template $\mathscr{L}(0,2v)$ if $v\leq 0$.
Proof.
It follows from Theorems 4.5 and 4.12. ∎
Corollary 4.14.
Let $L$ be the generalised T-link
$T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}),d)$
with $s_{k+1}\geqslant nr_{k+1}$. Then, $L$ is the $T^{2n}$-link
$T^{2n}((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}-nr_{k+1}),(r_{k+1};d)).$
Proof.
It follows from the proof of Theorem 4.12. ∎
Theorem 4.15.
Fix any negative integer $n$. Then, the generalised T-links
$T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}),d)$
with $s_{k+1}\geqslant nr_{k+1}$ consist of all links in $\mathbb{S}^{3}$.
Proof.
By Theorem 4.13, if we consider that $s_{k+1}\geqslant nr_{k+1}$, then the generalised T-links $T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}),d)$ are the set of all links in the Lorenz-like template $\mathscr{L}(0,2n)$. The Lorenz-like template $\mathscr{L}(0,2n)$ contains all links in $\mathbb{S}^{3}$ by Proposition 2.2. Therefore, the generalised T-links $T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}),d)$ with $s_{k+1}\geqslant nr_{k+1}$ consist of all links in $\mathbb{S}^{3}$. ∎
Corollary 4.16.
Generalised T-links $T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}),d)$ consist of all links in $\mathbb{S}^{3}$.
Proof.
It follows from Theorem 4.15 because, in particular, we can choose a negative integer $n$ and consider the stricter condition $s_{k+1}\geqslant nr_{k+1}$. ∎
Corollary 4.17.
Any link in the 3-sphere is the union of a $T$-link in [6, Definition 1.2] and a finite number of unknots. The upper bound in [6, Theorem 1.7] applies to the sum of the volumes of the hyperbolic pieces of all links in the 3-sphere.
Proof.
It follows from Corollary 4.16. ∎
Theorem 4.18.
Consider $L=T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}),d)$ a generalised T-link with $s_{k+1}\geqslant nr_{k+1}$ for some integer $n$. Let $\mathrm{Vol}(L)$ denote the sum of the volumes of the hyperbolic pieces of $\mathbb{S}^{3}-L$. Furthermore, let $\beta$ be the braid index of the Lorenz link
$L^{\prime}=T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}-nr_{k+1})).$
Then,
$\mathrm{Vol}(L)\leq 12v_{tet}(\beta^{2}+3\beta).$
Furthermore, this provides an upper bound for the volume of all hyperbolic pieces of each link complement in the 3-sphere.
Proof.
By Corollary 4.14, the generalised T-link $L$ is the $T^{2n}$-link
$L^{\prime\prime}=T^{2n}((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}-nr_{k+1}),(r_{k+1};d)).$
By Theorem 4.9, the sum of the volumes of the hyperbolic pieces of $\mathbb{S}^{3}-L^{\prime\prime}$ is less than or equal to $12v_{tet}(\beta^{2}+3\beta)$, where $\beta$ the braid index of the Lorenz link $L^{\prime}=T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}-nr_{k+1})).$ Therefore,
$\mathrm{Vol}(L)\leq 12v_{tet}(\beta^{2}+3\beta).$
Finally, from Corollary 4.16, this provides an upper bound for the volumes of the hyperbolic pieces of all links in the 3-sphere. ∎
In Theorem 4.18, we can find $n$ such that $0\leq s_{k+1}-nr_{k+1}<r_{k+1}$. Hence the braid index of the associated Lorenz links
$T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}-nr_{k+1}))$
can be less than $r_{k+1}$. Moreover, if $0\leq s_{k+1}-nr_{k+1}<r_{k+1}$, then braid index of $T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}-nr_{k+1}))$ can be even smaller than $r_{k}$. For example, the braid index of the Lorenz link $T(((4,3),(5,2),(7,2))$ is equal to 4, see [2, Corollary 8], [8, Section 3], or [9, Theorem 3.9].
When we compare the upper bounds for the volume obtained in Theorem 4.18 and [6, Theorem 1.7], we see that they provide better bounds for different set of parameters. For example, if the generalised T-link $L=T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}))$ has big $r_{k}$ but the Lorenz link $L^{\prime}=T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}-nr_{k+1}))$ has small braid index, then Theorem 4.18 provides a better upper bound for the volume of $L$ than [6, Theorem 1.7]. For example, consider $c>b>4$ and $a>0$. The generalised T-link $T((4,a),(b,2),(c,2+c))$ can have the number $b$ as big as we want but the braid index of the Lorenz link $T((4,a),(b,2),(c,2))$ is equal to $4$. On the other hand, if $r_{k}$ is less than 12 and the braid index of $L^{\prime}$ is $r_{k}$, then [6, Theorem 1.7] provides a better upper bound for the volume of $L$ than Theorem 4.18.
We ultimately derive an upper bound for the volumes of the hyperbolic pieces of all closed, orientable 3-manifolds based on the braid indexes of the Lorenz links.
Theorem 4.19.
For any closed, orientable and connected $3$-manifold $M$, there exists a generalised T-link
$L\coloneqq T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}),d)$
with $s_{k+1}\geqslant nr_{k+1}$ for some integer $n$ such that $M$ can be obtained by Dehn filling the complement of $L$ in the $3$-sphere. Furthermore, let $\mathrm{Vol}(M)$ be the sum of the volumes of the hyperbolic pieces of $M$ and let $\beta$ be the braid index of the Lorenz link
$T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}-nr_{k+1})).$
We have
$\mathrm{Vol}(M)\leq 12v_{tet}(\beta^{2}+3\beta).$
Proof.
Let $M$ be a closed, orientable and connected 3-manifold that is obtained by Dehn filling the complement of the generalised T-link
$T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}),d)$
with $s_{k+1}\geqslant nr_{k+1}$ in the 3-sphere.
Let $\mathrm{Vol}(L)$ be the sum of the volumes of the hyperbolic pieces of $\mathbb{S}^{3}-L$. By Theorem 4.18,
$\mathrm{Vol}(L)\leq 12v_{tet}(\beta^{2}+3\beta),$
where $\beta$ is the braid index of the Lorenz link
$L^{\prime}=T((r_{1},s_{1}),\dots,(r_{k},s_{k}),(r_{k+1},s_{k+1}-nr_{k+1})).$
Consider $\mathrm{Vol}(M)$ the sum of the volumes of the hyperbolic pieces of $M$. Since Dehn filling doesn’t increase volume by Thurston [21, Proposition 6.5.2], we have
$\mathrm{Vol}(M)\leq\mathrm{Vol}(L)\leq 12v_{tet}(\beta^{2}+3\beta).$
Finally, by the fundamental theorem of Wallace and Lickorish, all closed, orientable and connected 3-manifolds are obtained through Dehn filling of the complements of links in the 3-sphere [24, 19]. ∎
We note that the upper bound for the volume in [6, Theorem 1.7] can also be applied to all closed, orientable 3-manifolds if we consider Corollary 4.17.